Two years ago a man was 7 times older than his son but in three years he will be only four times as old as his son.How old are they?
Answer by Theo(13342)

m = number of years old the man is now.
s = number of years old his sone is now.

two years ago, the man was 7 times older than his son.

there is a difference between 7 times as old and 7 times older.
7 times as old says that a = 7 * b
7 times older says that a = b + 7*b = 8*b

your problem statement seems to mix them both.
i suspect you meant as old as, rather than older than.

i'll assume that, therefore, i'll rephrase the problem statement to say:

Two years ago a man was 7 times as old as his son, but in three years he will be only four times as old as his son.
How old are they?

your formulas are:

(m - 2) = 7 * (s - 2)
(m + 3) = 4 * (s + 3)

simplify these formulas to get:

m - 2 = 7 * s - 14
m + 3 = 4 * s + 12

solve for m in both equations to get:

m = 7 * s - 12
m = 4 * s + 9

subtract the second equation from the first to get 0 = 3 * s - 21
add 21 to both sides of this equation to get 21 = 3 * s
solve for s to get s = 7

2 years ago, the son was 5 and the man was 7 * 5 = 35
that means the man is currently 37 years old and his son is 7.

3 years from now, the man will be 40 and his son will be 10.
40 / 10 = 4 which means that the man will be 4 times as old as his son.

looks like man is 37 years old and his son is 7.

two years ago, 35/5 = 7
3 years from now, 40/10 = 4
solution looks good.
solution is man is 37 years old and his son is 7.